OpenStudy (anonymous):

A game is being played using one die. If the die is rolled and shows 4, the player wins $25. If the die shows any number other than 4 the player wins nothing. If there is a charge of $5 to play game what is the expected value?

OpenStudy (anonymous):

it is the amount you win times the probability you win it

OpenStudy (anonymous):

you win with probability \(\frac{1}{6}\) and lose with probability \(\frac{5}{6}\) it is \[20\times \frac{1}{6}-5\times \frac{5}{6}\] the reason for \(20\) instead of \(25\) is that even if you win, you already payed $5 to play